OpenStudy (anonymous):

Use the right triangle to solve the triangle b=4, c=6 Find a, A, B

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay, let's find the sides first.. Do you know the Pythagorean Theorem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

How could we apply this in this equation?

OpenStudy (anonymous):

To find a

OpenStudy (anonymous):

Brillant! Could you set up an equation for a?

OpenStudy (anonymous):

a^2+4^2=6^2 a= sqrt 20

OpenStudy (anonymous):

Awesome! Now have you learned what SOHCAHTOA is?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Nice! So we could say sin(B) = b/c right?

OpenStudy (anonymous):

It's crucial you understand that part, do you understand it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay so sin(B) = 2/3 and do arcsin on both sides with your calculator and you should get what B is!

OpenStudy (anonymous):

sin(B).. arcsin (sin(B)) = B

OpenStudy (anonymous):

I have to go but I hope you understand the jist of this.